# Bayesian Perspective for Orientation Estimation in Cryo-EM and Cryo-ET

### v1

@@ 2.2 The MMSE and MAP orientation estimators
Let $\hat{g}_{\mathrm{MAP}}$ be the MAP rotation estimator as defined in (2.10), and assume it is unique. Let $\hat{g}_{\mathrm{MMSE}}$ be the MMSE estimator as defined in (2.11), where $\Lambda$ is a uniform measure over $\mathsf{SO}(3)$. Then,
$\lim_{\sigma\to 0}\hat{g}_{\mathrm{MMSE}}=\hat{g}_{\mathrm{MAP}}.$ (2.13)
The proof of the proposition is provided in Appendix A. The proof relies on the existing result, [41, Theorem 5.10], and is valid not only to rotations but also to other group operators, such as translations.
@@ 3 Numerical methods for MMSE orientation estimation
This section compares the numerical performance of the MMSE and MAP estimators. We introduce various types of prior distributions for $\mathsf{SO}(3)$ beyond the uniform distribution and demonstrate how incorporating this prior knowledge can significantly improve the performance of the MMSE rotation estimator. Furthermore, we study the influence of the number of sampling points of the group $\mathsf{SO}(3)$ of 3D rotations (namely, the number of candidate rotations), denoted as $L$, on the quality of rotation estimation. For better illustration, we consider the simplified setting of model (2.3), where we assume $\tau_{i}^{2}=0$ for all $i$, which means $\tilde{\Sigma}=\sigma^{2}I_{d\times d}$ in (2.9).
Numerical procedure and the (possibly non-uniform) sampling of $\mathsf{SO}(3)$.
Generally, the denominator expression in (2.9), which involves the integral over $\mathsf{SO}(3)$, cannot be computed analytically and thus should be approximated numerically. To this end, we use Monte Carlo numerical approximation by sampling $\mathsf{SO}(3)$ with $L$ i.i.d. points $\left\{g^{(\ell)}\right\}_{\ell=0}^{L-1}\in\mathsf{SO}(3)$ according to the law $\Lambda$. Several approaches are available to generate samples following a general distribution $\Lambda$, such as the inverse transform sampling method and the rejection sampling method [20]. The method employed in this work is based on the inverse sampling theorem, as detailed in Appendix B. We mention that it is common to apply clever adaptive sampling procedures for acceleration in practice, see e.g., [44, 40].
Let us first consider the simplest scenario, where $\Lambda$ is a uniform distribution (i.e., the Haar measure on SO(3)). Denote $x_{\ell}:=\Pi\left((g^{(\ell)})^{-1}\cdot\overline{V}\right)$ for every $10\leq\ell\leq L-$ 1, where $\left\{g^{(\ell)}\right\}_{\ell=0}^{L-1}$ are generated i.i.d. according to the uniform distribution. Then, for every $10\leq\ell\leq L-$ 1, the posterior distribution for each rotation $g^{(\ell)}$ can be approximated by
$p^{(\ell)}:={\mathbb{P}}(g^{(\ell)}|y,\overline{V})\approx\frac{\exp\Big{(}-\left\|y-x_{\ell}\right\|^{2}/2\sigma^{2}\Big{)}}{\sum_{\ell=0}^{L-1}\exp\Big{(}-\left\|y-x_{\ell}\right\|^{2}/2\sigma^{2}\Big{)}},$ (3.1)
and the approximation improves as $L$ increases. The MMSE estimator, i.e., the Bayes estimator with respect to the Chordal distance, can then be approximated by
@@
$\hat{g}_{\mathsf{MMSE}}=\mathbb{E}[g|y,\overline{V}]\approx\sum_{\ell=0}^{L-1}g^{(\ell)}p^{(\ell)}=\frac{\sum_{\ell=0}^{L-1}g^{(\ell)}\exp\Big{(}-\left\|y-x_{\ell}\right\|^{2}/2\sigma^{2}\Big{)}}{\sum_{\ell=0}^{L-1}\exp\Big{(}-\left\|y-x_{\ell}\right\|^{2}/2\sigma^{2}\Big{)}}.$ (3.2)
To obtain $\hat{g}_{\mathsf{MMSE}}$ for a general non-uniform distribution, we only need to make very simple adjustments: the formula of numerical approximation (3.2) remains unchanged, and we simply need to draw samples $\left\{g^{(\ell)}\right\}_{\ell=0}^{L-1}$ according to the desired non-uniform rotational distribution over $\mathsf{SO}(3)$.
@@
Similarly, the MAP estimator in (2.10) relies on sampling or discretization for approximation, which can be approximated through grid search using the uniform rotation samples or pre-defined grid $\left\{g^{(\ell)}\right\}_{\ell=0}^{L-1}$ as
$\hat{g}_{\mathsf{MAP}}=$ $\underset{0\leq\ell\leq L-1}{\operatorname*{arg\,min}}\left\|y-x_{\ell}\right\|^{2}.$ (3.3)
Since both estimators require computing $\|y-x_{\ell}\|$ over all candidate rotations, each with a complexity of $O(d)$, the overall complexity of obtaining both estimators is at the same scale of $O(Ld)$. In many cases, it is preferable to work in a transformed basis (where rotating the volume yields greater accuracy and reduces the effective $d$), though this may introduce an additional $\log d$ factor [25].
@@
    $\hat{g}_{\mathsf{MMSE}}=\frac{\sum_{\ell=0}^{L-1}g^{(\ell)}\exp\Big{(}-\left\|y-\Pi\left(\left(g^{(\ell)}\right)^{-1}\cdot\bar{V}\right)\right\|^{2}/2\sigma^{2}\Big{)}}{\sum_{r=0}^{L-1}\exp\Big{(}-\left\|y-\Pi\left(\left(g^{(r)}\right)^{-1}\cdot\bar{V}\right)\right\|^{2}/2\sigma^{2}\Big{)}}.$ (3.4)
@@
    Perform the Orthogonal Procrustes algorithm (2.12) on $\hat{g}_{\mathsf{MMSE}}$ in (3.4) to get a proper rotation operator.
@@
One key advantage of the Bayesian framework is its flexibility in incorporating different prior distributions for the rotations. In practical cryo-EM applications, the distribution of rotations is often non-uniform [53, 30]. If the rotation distribution can be well estimated in early stage, the information of rotation distribution can then be integrated into the Bayes estimator to improve rotation estimation.
@@
For the uniform rotational distribution as discussed in the last section, the numerical MMSE estimator is computed using Monte Carlo sampling of the term in (2.11), under the assumption that $\tilde{\Sigma}=\sigma^{2}I_{d\times d}$ in model (2.8). To illustrate rotation estimation under different prior distributions on $\mathsf{SO}(3)$, we replace the uniform distribution on $\mathsf{SO}(3)$ with an isotropic Gaussian (IG) distribution on $\mathsf{SO}(3)$, denoted by $g\sim\mathcal{IG}_{\mathsf{SO}(3)}(\eta)$, parameterized by a scalar variance $\eta^{2}$. This distribution is frequently used in machine learning probabilistic models on $\mathsf{SO}(3)$ [10, 26, 22]. It is worth noting that the IG distribution serves as a typical example, and similar phenomena as observed here extend to other non-uniform distributions on $\mathsf{SO}(3)$ as well.
@@
The IG distribution $\mathcal{I}\mathcal{G}_{\mathsf{SO}(3)}(\eta)$ can be represented in an axis-angle form, with uniformly sampled axes of rotation and a rotation angle $\omega\in[0,\pi]$. The scalar variance $\eta^{2}$ controls the distribution of the rotation angle $\omega$: as $\eta\to\infty$, the $\mathsf{SO}(3)$ distribution approaches uniformity, whereas as $\eta\to 0$, $\omega$ becomes increasingly concentrated around 0, i.e., the rotation angle around the rotation axis is small. We apply the inverse sampling method to obtain i.i.d. samples from the IG distribution. Further details on this distribution are provided in Appendix B.
@@
Figure 2 illustrates how incorporating a prior distribution over $\mathsf{SO}(3)$ rotations, governed by the variance parameter $\eta$ of an isotropic Gaussian distribution affects, the accuracy of the MMSE rotation estimator. In all cases, the true underlying rotational distribution is modeled as an isotropic Gaussian distribution $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$ over $\mathsf{SO}(3)$. The MAP estimator was computed according to (3.3). For the MMSE estimators, the estimation process used different prior distributions with variance parameters $\eta=0.7,0.5$, and $0.1$, respectively. Specifically, the candidate rotations $g^{(\ell)}$ were generated according to these different priors (see (3.2)), highlighting the impact of prior mismatch on estimation accuracy. As the variance decreases (indicating a more concentrated and less uniform distribution closer to the true underlying distribution), the performance of the MMSE estimator improves, particularly at lower SNR conditions. These findings highlight the value of incorporating prior knowledge in rotation estimation, demonstrating its potential to substantially enhance accuracy. In stark contrast, the MAP estimator remains entirely unchanged regardless of the true underlying rotation distribution. As a result, its performance remains suboptimal, particularly in scenarios where the true distribution deviates significantly from uniformity.
@@
Figure 2: Impact of incorporating the prior rotational distribution on estimation accuracy. In all cases, the true underlying rotational distribution is modeled as an isotropic Gaussian distribution $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$. The metric used for comparison is the geodesic distance, as defined in (2.5). Here, $g$ denotes the true rotation, $\hat{g}_{\mathrm{MAP}}$ represents the MAP estimator from (3.3), and $\hat{g}_{\mathrm{MMSE}}$ denotes the Bayesian MMSE estimator from (3.2). The MMSE estimators are evaluated using different isotropic Gaussian distributions over the rotation group $\mathsf{SO}(3)$ with varying variances $\eta=0.7,0.5$, and $0.1$, respectively (see Appendix B). As the variance parameter $\eta$ decreases, resulting in less uniform rotational distribution closer to the true underlying rotational distribution, the accuracy of the MMSE rotation estimation improves. Each point in the curve plot is based on Monte Carlo simulations with 10,000 trials, using a sampling grid size of $L=300$. Below the curve plot, an illustration of colored 3D volumes, generating according to different rotational estimators, compares the performance of the MAP and MMSE estimators, where the second row on the right corresponds to using the true prior distribution $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$ in the MMSE estimation process. Clearly, the MMSE estimator leads to a more accurate rotation estimation.
@@
To study these effects, we restrict ourselves to the uniform rotational distribution setting. Figure 3 illustrates the impact of the sampling grid size of $\mathsf{SO}(3)$ together with different levels of SNR, on the geodesic distance between the MAP and MMSE rotation estimators relative to the true rotations. Several observations can be made. First, at high SNR levels, the MAP and MMSE estimators nearly coincide, as predicted by Proposition 2.1. In this regime, the geodesic distance scales empirically as $L^{1/3}$. This relationship can be attributed to the representation of the $\mathsf{SO}(3)$ rotation using three parameters, which leads to a typical grid size that scales as $L^{1/3}$ in each direction. As the SNR decreases, the disparity between the different estimators diminishes as a function of the grid sampling size $L$. Additionally, in the extreme case where $\sigma\to\infty$, the mean geodesic distance appears to be similar for both the MAP and MMSE estimators, showing no dependence on the grid size $L$.
@@
Figure 3: Impact of the sampling grid size of $\mathsf{SO}(3)$ ($L$) and the signal-to-noise ratio (SNR) on rotation estimation accuracy. This figure shows the accuracy of rotation estimation under varying sampling grid sizes $L$ of the rotation group $\mathsf{SO}(3)$ and different SNR levels of the observed data $y$ in the model (2.1). The metric used for comparison is the geodesic distance, as defined in (2.5). Here, $g$ denotes the true rotation, $\hat{g}_{\mathrm{MAP}}$ represents the MAP estimator from (3.3), and $\hat{g}_{\mathrm{MMSE}}$ denotes the Bayesian MMSE estimator from (3.2). In the high SNR regime ($\sigma\to 0$), the MAP and MMSE estimators converge, and the geodesic distance scales empirically as $\propto L^{1/3}$. This scaling reflects the three-parameter nature of $\mathsf{SO}(3)$ rotations, where the resolution of the sampling grid improves as $L$ increases. The results shown are based on Monte Carlo simulations with 3000 trials per data point.
4 Bayesian rotation determination as part of the volume reconstruction problem

### v2

@@ 2.3 The MLE, MAP and MMSE orientation estimators
The Bayes estimator of the loss function defined by the chordal distance $\hat{g}_{\mathsf{MMSE}}$ (2.11) is equal to the Orthogonal Procrustes solution (2.13) applied on the intermediate estimator $\hat{g}_{\mathsf{relax}}$, i.e.,
$\hat{g}_{\mathsf{MMSE}}={\operatorname*{arg\,min}}_{\Omega\in\mathsf{SO}(3)}\ \left\|\hat{g}_{\mathsf{relax}}-\Omega\right\|_{F}.$ (2.14)
The MMSE, MAP and MLE estimators in the high SNR regime.
We conclude the section by showing the next proposition. It shows that the MMSE and MAP estimators converge in the high SNR regime (i.e., $\sigma\to 0$), as demonstrated empirically in Figures 3, 4, and 5. This implies that the MMSE estimator’s superior statistical properties are most advantageous in low SNR conditions, which are common in structural biology applications like cryo-EM and cryo-ET bendory2020single . In these low SNR environments, the MMSE estimator consistently outperforms its MAP counterpart.
Proposition 2.2.
Let $\Lambda$ be any distribution over $\mathsf{SO}(3)$ with a strictly positive density, bounded below by some constant $c>0$. Let $\hat{g}_{\mathsf{MLE}}$, $\hat{g}_{\mathsf{MAP}}$ be the MLE and MAP rotation estimators as defined in (2.9) and (2.10), respectively, and assume they are unique. Let $\hat{g}_{\mathsf{MMSE}}$ be the MMSE estimator as defined in (2.11). Then, we have,
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MMSE}}=\lim_{\sigma\to 0}\hat{g}_{\mathsf{MAP}}=\hat{g}_{\mathsf{MLE}}.$ (2.15)
The proof of the proposition is provided in Appendix D. The proof relies on the existing result, (robert1999monte, , Theorem 5.10), and is valid not only to rotations but also to other group operators, such as translations.
@@ 3 Numerical methods for MMSE orientation estimation
This section compares the numerical performance of the MMSE and MLE estimators. We also introduce various types of prior distributions for $\mathsf{SO}(3)$ beyond the uniform distribution and demonstrate how incorporating this prior knowledge can significantly improve the performance of the MMSE rotation estimator. Furthermore, we study the influence of the number of sampling points of the group $\mathsf{SO}(3)$ of 3D rotations (namely, the number of candidate rotations), denoted as $L$, on the quality of rotation estimation. For better illustration, we consider the simplified setting of model (2.1), where $\Sigma=\sigma^{2}I_{d\times d}$, although our method applies to the general covariance setting as well.
Numerical procedure and sampling of $\mathsf{SO}(3)$.
Since the posterior distribution of $g\in\mathsf{SO}(3)$ in (2.6) is continuous, a numerical discretization of the rotation group is required. In this work, the expectation is approximated using a quadrature rule over $\mathsf{SO}(3)$. Let $\{(g^{(\ell)},w_{\ell})\}_{\ell=0}^{L-1}$ denote a numerical quadrature on $\mathsf{SO}(3)$, constructed as a product of a spherical quadrature on $\mathsf{S}^{2}$ and a circular rule over the Euler angles graf2009sampling ; graf2012unified . The implementation employed here follows the publicly available MATLAB code described in hoskins2024subspace .
We first consider the uniform case, where $\Lambda$ is the Haar distribution on $\mathsf{SO}(3)$. For each quadrature node, define $x_{\ell}:=\Pi\!\big(g^{(\ell)}\circ{V}\big)$, for $0\leq\ell\leq L-1$, where $\{g^{(\ell)}\}_{\ell=0}^{L-1}$ are the quadrature nodes and $\{w_{\ell}\}_{\ell=0}^{L-1}$ are their associated weights. The MMSE estimator can then be expressed as the weighted average
@@
$\hat{g}_{\mathsf{MMSE}}=\mathbb{E}[g\mid y,{V}]$ $\approx\frac{\displaystyle\sum_{\ell=0}^{L-1}g^{(\ell)}\,w_{\ell}\exp\!\Big(-\tfrac{\|y-x_{\ell}\|^{2}}{2\sigma^{2}}\Big)}{\displaystyle\sum_{\ell=0}^{L-1}w_{\ell}\exp\!\Big(-\tfrac{\|y-x_{\ell}\|^{2}}{2\sigma^{2}}\Big)}=\sum_{\ell=0}^{L-1}g^{(\ell)}\,p^{(\ell)},$ (3.1)
where the posterior quadrature weights are defined as
$p^{(\ell)}\;\triangleq\;\frac{w_{\ell}\exp\!\Big(-\tfrac{\|y-x_{\ell}\|^{2}}{2\sigma^{2}}\Big)}{\displaystyle\sum_{r=0}^{L-1}w_{r}\exp\!\Big(-\tfrac{\|y-x_{r}\|^{2}}{2\sigma^{2}}\Big)},\qquad 0\leq\ell\leq L-1.$ (3.2)
For a general non-uniform prior $\Lambda(g)$, the same quadrature can be applied with modified posterior weights:
$p^{(\ell)}\;\propto\;w_{\ell}\,\Lambda\!\big(g^{(\ell)}\big)\,\exp\!\Big(-\tfrac{\|y-x_{\ell}\|^{2}}{2\sigma^{2}}\Big),$ (3.3)
followed by normalization to ensure $\sum_{\ell=0}^{L-1}p^{(\ell)}=1$.
@@
Similarly, the MLE estimator in (2.9) relies on the discretization of $\mathsf{SO}(3)$, which can be approximated through grid search using the uniform rotation samples or pre-defined grid $\left\{g^{(\ell)}\right\}_{\ell=0}^{L-1}$ as
$\hat{g}_{\mathsf{MLE}}=$ $\underset{0\leq\ell\leq L-1}{\operatorname*{arg\,min}}\left\|y-x_{\ell}\right\|^{2}.$ (3.4)
Both estimators require evaluating $\|y-x_{\ell}\|$ for all candidate rotations, which naively costs $O(d)$ per candidate and thus $O(Ld)$ overall. This estimate, however, overlooks the structure of $\mathsf{SO}(3)$. Any 3D rotation can be decomposed into a viewing direction (two degrees of freedom, discretized with about $L^{2/3}$ samples) and an in-plane rotation about that direction (one degree of freedom, discretized with $L^{1/3}$ samples). By exploiting FFT-based methods, all $L^{1/3}$ in-plane rotations can be evaluated simultaneously for each viewing direction. Consequently, the effective complexity becomes $O(L^{2/3}\,d\,\log d)$, which improves over the naive bound by a factor of $L^{1/3}$ kostelec2008ffts ; kileel2024fast . This reduction is substantial in practice, since $L$ often reaches tens or even hundreds of thousands in high-resolution cryo-EM.
@@
    $\hat{g}_{\mathsf{relax}}=\frac{\sum_{\ell=0}^{L-1}g^{(\ell)}\exp\Big(-\left\|y-\Pi\left({g^{(\ell)}}\circ{V}\right)\right\|^{2}/2\sigma^{2}\Big)}{\sum_{\ell=0}^{L-1}\exp\Big(-\left\|y-\Pi\left({g^{(\ell)}}\circ{V}\right)\right\|^{2}/2\sigma^{2}\Big)}.$ (3.5)
@@
    Perform the Orthogonal Procrustes algorithm (2.13) on $\hat{g}_{\mathsf{relax}}$ in (3.5) to get the proper rotation estimator $\hat{g}_{\mathsf{MMSE}}$.
@@
One key advantage of the Bayesian framework is its flexibility in incorporating different prior distributions for the rotations. In practical cryo-EM applications, the distribution of particle orientations is often non-uniform due to preferred particle orientations or sample preparation artifacts tan2017addressing ; lyumkis2019challenges . If the rotation distribution can be well estimated in an early stage, the information of the rotation distribution can then be integrated into the Bayes estimator to improve rotation estimation.
@@
To illustrate rotation estimation under different prior distributions on $\mathsf{SO}(3)$, we replace the uniform distribution on $\mathsf{SO}(3)$ with an isotropic Gaussian (IG) distribution on $\mathsf{SO}(3)$, denoted by $g\sim\mathcal{IG}_{\mathsf{SO}(3)}(\eta)$, parameterized by a scalar variance $\eta^{2}$. This distribution is frequently used in machine learning probabilistic models on $\mathsf{SO}(3)$ corso2022diffdock ; leach2022denoising ; jagvaral2024unified . It is worth noting that the IG distribution serves as a typical example, and similar phenomena as observed here extend to other non-uniform distributions on $\mathsf{SO}(3)$ as well.
@@
The IG distribution $\mathcal{I}\mathcal{G}_{\mathsf{SO}(3)}(\eta)$ can be represented in an axis-angle form, with uniformly sampled axes of rotation and a rotation angle $\omega\in[0,\pi]$. The scalar variance $\eta^{2}$ controls the distribution of the rotation angle $\omega$: as $\eta\to\infty$, the $\mathsf{SO}(3)$ distribution approaches uniformity, whereas as $\eta\to 0$, $\omega$ becomes increasingly concentrated around 0, i.e., the rotation angle around the rotation axis is small. We apply the inverse sampling method to obtain i.i.d. samples from the IG distribution. Further details on this distribution are provided in Appendix E.
@@
Figure 2 illustrates how incorporating a prior distribution over $\mathsf{SO}(3)$ rotations, governed by the variance parameter $\eta$ of an isotropic Gaussian distribution affects, the accuracy of the MMSE rotation estimator. In all cases, the true underlying rotation distribution is modeled as an isotropic Gaussian distribution $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$ over $\mathsf{SO}(3)$. The MLE estimator was computed according to (3.4). For the MMSE estimators, the estimation process used different prior distributions with variance parameters $\eta=0.7,0.5$, and $0.1$, respectively. Specifically, the candidate rotations $g^{(\ell)}$ were generated according to these different priors (see (3.1)), highlighting the impact of prior mismatch on estimation accuracy. As the variance decreases (indicating a more concentrated and less uniform distribution closer to the true underlying distribution), the performance of the MMSE estimator improves, particularly at lower SNR conditions. These findings highlight the value of incorporating prior knowledge in rotation estimation, demonstrating its potential to enhance accuracy substantially. In stark contrast, the MLE estimator remains entirely unchanged regardless of the true underlying rotation distribution. As a result, its performance remains suboptimal, particularly in scenarios where the true distribution deviates significantly from uniformity.
@@
Figure 2: Impact of incorporating the prior rotation distribution on estimation accuracy. Simulations are performed for the cryo-ET model (1.2), excluding the projection step. The true rotation distribution is modeled as an isotropic Gaussian $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$. Estimation performance is measured using the geodesic distance defined in (2.8). Here, $g$ denotes the true rotation, $\hat{g}_{\mathrm{MLE}}$ is the maximum-likelihood estimator from (3.4), $\hat{g}_{\mathrm{MAP}}$ is the maximum a posteriori estimator from (2.10), and $\hat{g}_{\mathrm{MMSE}}$ denotes the Bayesian minimum mean square error estimator from (3.1). The MMSE estimators are computed assuming isotropic Gaussian priors on $\mathsf{SO}(3)$ with different concentration parameters $\eta\in\{0.5,0.1\}$ (see Appendix E). As $\eta$ decreases, the prior becomes more concentrated and closer to the true underlying distribution, leading to improved accuracy of both the MAP and MMSE estimators. Each data point in the plot is averaged over 3000 Monte Carlo trials using a rotation grid of size $L=2976$. Below the curve plot, we show visual reconstructions of denoised 3D volumes generated using different rotational estimators. The second row (right panel) corresponds to the MMSE estimator that employs the true prior $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$. Incorporating this prior clearly enhances rotational accuracy compared to the MLE baseline.
@@
To study these effects, we restrict ourselves to the uniform rotation distribution setting. Figure 3 illustrates the impact of the sampling grid size of $\mathsf{SO}(3)$ together with different levels of SNR, on the geodesic distance between the MLE and MMSE rotation estimators relative to the true rotations. Several observations can be made. First, at high SNR, the MLE and MMSE estimators nearly coincide, consistent with Proposition 2.2. In this regime, the geodesic error decreases with the grid resolution, scaling empirically as $L^{1/3}$. This behavior arises because a discretization of $\mathsf{SO}(3)$ involves three angular parameters, so the resolution in each parameter direction grows like $L^{1/3}$. Second, as the SNR decreases, noise dominates and the advantage of refining the grid diminishes: both estimators approach similar performance, and the dependence on $L$ becomes weaker. In the extreme limit $\sigma\to\infty$, the mean geodesic distance of the MLE and MMSE estimators is indistinguishable and shows no dependence on the grid size $L$.
@@
Figure 3: Impact of the sampling grid size of $\mathsf{SO}(3)$ ($L$) and the signal-to-noise ratio (SNR) on rotation estimation accuracy. This figure shows the accuracy of rotation estimation under varying sampling grid sizes $L$ of the rotation group $\mathsf{SO}(3)$ and different SNR levels of the observed data $y$ in the model (2.1). Simulations are performed for the cryo-ET model (1.2), excluding the projection step. The metric used for comparison is the geodesic distance, as defined in (2.8). Here, $g$ denotes the true rotation, $\hat{g}_{\mathrm{MLE}}$ represents the MLE estimator from (3.4), and $\hat{g}_{\mathrm{MMSE}}$ denotes the Bayesian MMSE estimator from (3.1). In the high SNR regime ($\sigma\to 0$), the MLE and MMSE estimators converge, and the geodesic distance scales empirically as $\propto L^{1/3}$. This scaling reflects the three-parameter nature of $\mathsf{SO}(3)$ rotations, where the resolution of the sampling grid improves as $L$ increases. The results shown are based on Monte Carlo simulations with 3000 trials per data point.
4 Bayesian orientation estimation as part of the volume reconstruction problem